OpenStudy (anonymous):

What is the quotient: (56x^3 + 16x^2 – 32) ÷ 8 ?

OpenStudy (anonymous):

factor out an 8 from the divisor (numerator)... that'll cancel with the denominator (divisor)

OpenStudy (anonymous):

***dividend....

OpenStudy (anonymous):

Sorry, the "Equation" button is not responding with the Firefox browser.

OpenStudy (anonymous):

write it out? :)

OpenStudy (anonymous):

7 x^3 + 2 x^2 - 4 is the result of dividing each term by 8 as recommended by @dpainc

Parth (parthkohli):

Factoring, you get \(\Large \color{MidnightBlue}{\Rightarrow 8(7x^3 + 2x^2 - 4) = 56x^3 + 16x^2 - 32 }\)

Parth (parthkohli):

You just factor out the 8 and see the terms you have to multiply to it for getting the dividend.

OpenStudy (anonymous):

From the Chrome browser: \[\frac{\left(8 \left(7 x^3+2 x^2-4\right)\right)}{8}=7 x^3+2 x^2-4 \]

OpenStudy (anonymous):

Thank you for the medal, @dpalnc